OpenStudy (anonymous):

Find the sum. 2 3/5 + 1 3/7=

hero (hero):

Can't use decimals on this one because of the 3/7

hero (hero):

But... 13/5 + 10/7 we can add that

OpenStudy (anonymous):

Yup but i get a decimal and it has to be a fraction and I get 3.028571

hero (hero):

You can't use decimals for this one :P

OpenStudy (anonymous):

I knowww D:

hero (hero):

2 3/5 + 1 3/7 = (2+1)(3/5 + 3/7) = 3(21/35 + 9/35) = 3 30/35 = 3 6/7

hero (hero):

short cut way to do it :P

OpenStudy (anonymous):

Can not be a 3 at the front unless it's 3 1/2

hero (hero):

oh wait... I made a mistake..

OpenStudy (anonymous):

ik it's not! 4 1/35 or 4 4/35

OpenStudy (anonymous):

so it must be 4 1/7

OpenStudy (anonymous):

orr 3 1/2 :D

hero (hero):

2 3/5 + 1 3/7 = (2+1)(3/5 + 3/7) = 3(21/35 + 15/35) = 3 36/35 = 4 1/35

hero (hero):

I'm convinced you must have wrote something wrong now

OpenStudy (anonymous):

how is it 4 1/35! D:

OpenStudy (anonymous):

i did i did Lol Yup i caught my mistake :D

hero (hero):

Okay, re-post the correct one

OpenStudy (anonymous):

oh my question is correct but when I worked it out i was wrong! D:

OpenStudy (anonymous):

but no i fixed it the answer is B. a 1 1/35 i believe

OpenStudy (anonymous):

2 3/5 + 1 3/7 = (2+1) (3/5 + 3/7) = 3(21/35 + 15/35) = 3 36/35 = 4 1/35

hero (hero):

That's impossible for the answer to be 1 1/35

OpenStudy (anonymous):

yeah it's 4 1/35

OpenStudy (anonymous):

you said 1 and it's 4

OpenStudy (anonymous):

well i made a typing mistake up there lol i meant 4

OpenStudy (anonymous):

*ment

OpenStudy (anonymous):

the answer is B. a 4 1/35 i believe

hero (hero):

My answer is correct. I said it was 4 1/35

OpenStudy (anonymous):

:O lol

hero (hero):

All you have to do is scoll up. You can see where I posted my answer.

OpenStudy (anonymous):

Oh yeah! haha ok

hero (hero):

I don't know why you get so confused on these

OpenStudy (anonymous):

Hero help find the difference on the next question

hero (hero):

just a sec lunch